THE C_{3}−BENDING VIBRATIONAL LEVELS OF C_{3} −NE AND C_{3} −XE VAN DER WAALS COMPLEXES

JUN-MEI CHAO, GUIQIU ZHANG, AND \underline{YEN-CHU HSU}, INSTITUTE OF ATOMIC AND MOLECULAR SCIENCES, ACADEMIA SINICA, P. O. BOX 23-166, TAIPEI 107, TAIWAN, R. O. C..

The C_{3} − Ne and C_{3} − Xe van der Waals (vdW) complexes were produced from 193-nm photolysis of an allene-rare gas atom gas mixture under supersonic expansion. The C_{3}−bending vibrational levels (v_{b}=2-10) of the ground electronic states of these two complexes have been obtained from the wavelength-resolved emission from vdW levels observed near the C_{3}, Ã-X̃, 2_{0}^{2−}K_{0}^{1}, 2_{0}^{4−}K_{0}^{1}, 2_{0}^{2+}K_{0}^{1}, and 1_{0}^{1}K_{0}^{1} bands. A ”free rotor” model was used to fit the observed C_{3}−bending vibrational levels to parameters of dipole-induced dipole interaction of these complexes. Comparison of level structure observed among four C_{3} − Rg complexes will be presented.